# AACR-Bench: Evaluating Automatic Code Review with Holistic Repository-Level Context

### v1

@@ 4.2.3 Languange-wise impact on ACR performance
Meanwhile, a cross-comparison of all mainstream frameworks (No context, BM25, Embedding, Agent) also reveals that C# exhibits exceptional interpretability and model-friendliness, whereas the C language consistently remains at the bottom of the performance spectrum. In the context-free “No context” mode, GPT-5.2 achieved the highest F1 score across the board (0.309) on C#, while the same model scored only 0.085 on C. This trend is further corroborated by Qwen-480B-Coder (C# $0.268$ vs. C $0.104$). Given the application status of both languages, we can reasonably assume that C# and C provide comparable abundance for training corpora. In this sense, this phenomenon deeply reflects the impact of programming language intrinsic characteristics on the performance of LLMs. For example, as a strongly typed, and object-oriented language, C# possesses rigorous namespace management and explicit type definitions. In contrast, C relies heavily on pointer operations, macro definitions, and implicit memory management. Its dependency information is often implicit within unstructured header files or linking logic. We hypothesize that such structural features exert a substantial influence on the performance of LLMs in ACR tasks.
@@
5 Error Case Analysis
Detailed case studies of hallucinated review comments generated by the model are provided in the Appendix.
@@ 6 Conclusion
We proposed AACR-Bench, a multi-lingual benchmark designed to evaluate ACR systems with repository-level context. AACR-Bench fills a critical gap in the field by better measuring the ability of various methods to leverage complex, repo-level context to perform ACR. Our extensive empirical evaluation on AACR-Bench reveals that the granularity/level of context and the choice of retrieval methods significantly impact ACR performance, and this influence varies depending on the LLM, programming language, and the LLM usage paradigm (e.g., whether an Agent architecture is employed). This finding strongly underscores the necessity of constructing a dataset such as AACR-Bench.
Like any benchmark, our work has limitations. Specifically, although we leveraged LLM-generated reviews to augment our dataset, constructing a fully comprehensive Ground Truth remains a formidable challenge due to the inherent complexity and subjectivity of real-world software systems. Future work will focus on further expanding the dataset scale and exploring more advanced semi-automated methods to refine the Ground Truth quality.
@@
Redefining Context Understanding
This work reframes the evaluation of “context effectiveness” in ACR. We argue that high-quality code review depends less on Information Retrieval and more on Information Utilization and reasoning robustness. Our findings indicate that without sufficient noise tolerance, even relevant code retrieved via BM25 or Embeddings can degrade performance. The superior reasoning of Agents in complex scenarios suggests a paradigm shift from “Passive Code Ingestion” (receiving pre-retrieved snippets) to “ctive Code Auditing”. Enabling models to actively explore context, verify hypotheses through multi-turn interactions, and filter noise—mimicking human expert behavior—represents the most promising direction for achieving reliable automated code review.
@@ B.1 Detailed Dataset Curation Process
  • Review Richness: The PR must contain at least two inline comments, including at least one constructive comment that was accepted (i.e., resulted in code modifications).
@@
Constrained by the cognitive boundaries and capacity limits of human reviewers, relying solely on manual efforts often fails to uncover all potential code defects. To address the resulting issue of ”Label Incompleteness,” we employed LLM generation techniques to comprehensively augment the review comments. To mitigate single-model bias and ensure diversity, we constructed a generation matrix comprising six mainstream open-source and proprietary models. These models generated review comments in parallel through two heterogeneous frameworks: an Internal Review System and an Open-Source Agent System (Claude Code). All generated comments underwent semantic de-duplication before being merged with the previously augmented human review comments, forming a candidate set for verification. This semantic de-duplication was performed using Qwen3-235B-A22B-Thinking-2507. Specifically, all review comments were first grouped based on their repository, Pull Request (PR), file path, and the specific Diff Hunk addressed. Within each group, comments underwent pairwise comparison via the LLM, and duplicates were removed based on the comparison results. The prompt used for this comparison is shown in Figure 9. We use the election of results running 5 times of judgement.
@@
Determine whether two given review comments express the same concern or suggestion regarding the provided diff hunk. Ignore differeneces in wording, tone, or formatting-focus solely on semantic equivalence of the underlying issue. If the core intent and technical substance are identical, answer "yes"; otherwise, answer "no".
@@ C.1 Detailed Benchmark Process
The evaluation process of AACR-Bench aligns with the Diff-oriented paradigm of modern code review. Within the benchmark, each Pull Request (PR) serves as an evaluation instance, where the collection of review comments associated with that PR constitutes the Ground Truth. For each PR, the Automated Code Review (ACR) method under evaluation is required to iterate through the Diff Hunks within the code changes and generate review comments for each hunk. The code review capability of the ACR method is then measured by comparing the generated comments against the Ground Truth to calculate Precision, Recall, and F1-score. All the models are run with $Temperature=0.7$, $Top_{p}=0.95$, $Top_{k}=20$.
@@ 3. Include a copyable prompt for Claude Code that the user can use to fix the issue, formatted as:
Please conduct a code review for the changes from {base_commit}...{target_commit}, **write the review comments to comments.txt(if the file exists, clear it first)**,
@@ Appendix D Case Study
In this section, we present case studies illustrating correct and incorrect review comments generated by models under Agent-based and similarity-based retrieval settings. For each case, we provide the Repo, PR ID, File Path, along with the Diff Hunk in the file under review, the generated review comment, the retrieved code context (where applicable), and an analysis of the erroneous comments. We observe that current models continue to suffer from knowledge hallucinations during code review. Furthermore, noisy data introduced by context retrieval is identified as a significant factor contributing to the generation of incorrect review comments.

### v2

@@ 4.2.3 Languange-wise impact on ACR performance
Meanwhile, a cross-comparison of all mainstream frameworks (No context, BM25, Embedding, Agent) also reveals that C# exhibits exceptional interpretability and model-friendliness, whereas the C language consistently remains at the bottom of the performance spectrum. In the context-free “No context” mode, GPT-5.2 achieved the highest F1 score across the board (0.309) on C#, while the same model scored only 0.085 on C. This trend is further corroborated by Qwen-480B-Coder (C# $0.268$ vs. C $0.104$). Given the application status of both languages, we can reasonably assume that C# and C provide comparable abundance for training corpora. In this sense, this phenomenon deeply reflects the impact of programming language intrinsic characteristics on the performance of LLMs. For example, as a strongly typed, and object-oriented language, C# possesses rigorous namespace management and explicit type definitions. In contrast, C relies heavily on pointer operations, macro definitions, and implicit memory management. Its dependency information is often implicit within unstructured header files or linking logic. We hypothesize that such structural features exert a substantial influence on the ACR performance of LLMs.
@@
5 Error Analysis
Detailed case studies of incorrect review comments generated by the model are provided in the Appendix D. We observe that current models continue to suffer from knowledge errors during code review. Furthermore, noisy data introduced by context retrieval is identified as a significant factor contributing to the generation of incorrect review comments.
@@ 6 Conclusion
We proposed AACR-Bench, a multi-lingual benchmark designed to evaluate ACR systems with repository-level context. AACR-Bench fills a critical gap in the field by better measuring the ability of various methods to leverage complex, repo-level context to perform ACR. Our extensive empirical evaluation on AACR-Bench reveals that:
The granularity/level of context and the choice of retrieval methods significantly impact ACR performance, and this influence varies depending on the LLM, programming language, and the LLM usage paradigm (e.g., whether an Agent architecture is employed).
As long as the above finding strongly underscores the necessity of constructing a dataset such asAACR-Bench, our work also has limitations. Specifically, although we leveraged LLM-generated reviews to augment our dataset, constructing a fully comprehensive Ground Truth remains a formidable challenge due to the inherent complexity and subjectivity of real-world software systems. Future work will focus on further expanding the dataset scale and exploring more advanced semi-automated methods to refine the Ground Truth quality.
@@
From Passive Ingestion to Active Auditing
This work reframes the evaluation of “context effectiveness” in ACR. We argue that high-quality code review depends less on Information Retrieval and more on Information Utilization and reasoning robustness. Our findings indicate that without sufficient noise tolerance capability, even relevant code retrieved via BM25 or Embeddings can degrade performance. The superior reasoning of Agents in complex scenarios suggests a paradigm shift from “Passive Code Ingestion” (receiving pre-retrieved snippets) to “Active Code Auditing”. Enabling models to actively explore context, verify hypotheses through multi-turn interactions, and filter noise—mimicking human expert behavior—represents the most promising direction for achieving reliable automated code review.
@@ Appendix A Overview of the Appendix
The Appendix is organized as follows:
  • Section B provides detailed information on AACR-Bench, including its data construction process, comparisons with existing benchmarks, and representative data samples.
  • Section C describes the experimental setup of AACR-Bench, covering the full evaluation pipeline, tested models, hyperparameter settings, and an analysis of fine-grained results.
  • Section D provides case studies based on the AACR-Bench evaluation results
@@ B.1 Detailed Dataset Curation Process
  • Review Richness: The PR must contain at least two inline comments, including at least one constructive comment that was accepted (i.e., resulted in code modifications)..
@@
Constrained by the cognitive boundaries and capacity limits of human reviewers, relying solely on manual efforts often fails to uncover all potential code defects. To address the resulting issue of “Incomplete Issue Annotation,” we employed LLM generation techniques to comprehensively augment the review comments. To mitigate single-model bias and ensure diversity, we constructed a generation matrix comprising six mainstream open-source and proprietary models. These models generated review comments in parallel through two heterogeneous frameworks: an Internal Review System and an Open-Source Agent System (Claude Code). All generated comments underwent semantic de-duplication before being merged with the previously augmented human review comments, forming a candidate set for verification. This semantic de-duplication was performed using Qwen3-235B-A22B-Thinking-2507. Specifically, all review comments were first grouped based on their repository, Pull Request (PR), file path, and the specific Diff Hunk addressed. Within each group, comments underwent pairwise comparison via the LLM, and duplicates were removed based on the comparison results. The prompt used for this comparison is shown in Figure 9. We use the election of results running 5 times of judgement.
@@
Determine whether two given review comments express the same concern or suggestion regarding the privided diff hunk. Ignore differences in wording, tone, or formatting-focus solely on semantic equivalence of the underlying issue. If the core intent and technical substance are identical, answer "yes", otherwise, answer "no".
@@ C.1 Detailed Benchmark Process
The evaluation process of AACR-Bench aligns with the Diff-oriented paradigm of modern code review. Within the benchmark, each Pull Request (PR) serves as an evaluation instance, where the collection of review comments associated with that PR constitutes the Ground Truth. For each PR, the Automated Code Review (ACR) method under evaluation is required to iterate through the Diff Hunks within the code changes and generate review comments for each hunk. The code review capability of the ACR method is then measured by comparing the generated comments against the Ground Truth to calculate Precision, Recall, and F1-score. Prompt for review comment match is the same as 9 and the model used is Qwen3-235B-A22B-Instruct-2507. A generated review comment is considered Line Correct if its line range overlaps with that of a review comment in the ground truth. It is considered Semantically Correct if it overlaps with the ground truth in terms of line numbers and the content is identical. All the models are run with $Temperature=0.7$, $Top_{p}=0.95$, $Top_{k}=20$.
@@ 3. Include a copyable prompt for Claude Code that the user can use to fix the issue, formatted as:
Please conduct a code review for the changes from {base_commit}...{target_commit}, with PR title: {pr_title}
PR description: {pr_description}
**write the review comments to comments.txt(if the file exists, clear it first)**,
@@ C.2 Additional Statistics
Table 9 and Table 10 shows additional statistics in experiment.
@@ Appendix D Case Study
In this section, we present case studies illustrating correct and incorrect review comments generated by models under Agent-based and similarity-based retrieval settings. For each case, we provide the Repo, PR ID, File Path, along with the Diff Hunk in the file under review, the generated review comment, the retrieved code context (where applicable), and an analysis of the erroneous comments. We observe that current models continue to suffer from knowledge errors during code review. Furthermore, noisy data introduced by context retrieval is identified as a significant factor contributing to the generation of incorrect review comments.